Let f(x,y)∈R be a function of two variables. Then the gradient of f

is
Select one:
a. equal to the directional derivative of f

b. a vector in R2

c. a vector in R3


d. a scalar
1
Expert's answer
2020-05-25T21:13:54-0400

The gradient of a function f(x,y) in two variables x and y is the vector function in R2 given by:

"< \frac{ \delta f} { \delta x } (x,y) , \frac{ \delta f} { \delta y } (x,y) >"

For example, suppose f(x,y) = x3 + y2

then, the gradient of f(x,y) is

"< \frac{ \delta f} { \delta x } (x,y) , \frac{ \delta f} { \delta y } (x,y) >"


"=< \frac{ \delta (x^3 + y^2 )} { \delta x } , \frac{ \delta (x^3 + y^2 )} { \delta y }>"


"= < 3x^2 , 2y>"

This is a vector in R2

So, the correct answer is option B.